OpenStudy (anonymous):

need help please. studying for my final and need someone to remind me how to do these. for a linear function. f(4)=11 and f(-2)=2. find f(3). I know that the equation for a linear function is f(x)=mx+b

ganeshie8 (ganeshie8):

you have two points : (4, 11) adn (-2, 2) right ?

ganeshie8 (ganeshie8):

f(4) = 11 means you're sending in x=4, and getting the value y=11 back. that forms a point on the function

OpenStudy (anonymous):

so would i try to find m first?

OpenStudy (anonymous):

with the given points

ganeshie8 (ganeshie8):

you can do it in multiple ways. i prefer using the two points equation of line : equation of line passing through (x1, y1) and (x2, y2) : \(\huge y-y1 = \frac{y2-y1}{x2-x1}(x-x1) \)

ganeshie8 (ganeshie8):

equation of line passing thru (4, 11) adn (-2, 2) \(\huge y-11 = \frac{2-11}{-2-4}(x-4)\)

OpenStudy (anonymous):

for m i got 1.5

ganeshie8 (ganeshie8):

\(\huge y-11 = 1.5(x-4)\)

ganeshie8 (ganeshie8):

\(\huge y-11 = 1.5x-6\) \(\huge y = 1.5x + 5\)

ganeshie8 (ganeshie8):

to find f(3), plugin x=3 in above equation !

OpenStudy (anonymous):

so i only really need to use one set of the points? since i was given (4,11) and (-2,2)

ganeshie8 (ganeshie8):

it helps to know equations of lines in 3 forms : 1) slope, intercept form : y = mx+c 2) point, slope form : y-y1 = m(x-x1) 3) two points form : y-y1 = (y2-y1/x2-x1)(x-x1)

ganeshie8 (ganeshie8):

we need both points. we used both points to get the equation

OpenStudy (anonymous):

but in my case i would be able to use the first one right?

OpenStudy (anonymous):

so f(x)=9.5?

OpenStudy (anonymous):

*f(3)=9.5

ganeshie8 (ganeshie8):

thats right!!

OpenStudy (anonymous):

thank you!

ganeshie8 (ganeshie8):

yw..